# Blessing or curse? A survey on the Impact of Generative AI on Fake News

### v1

@@ I Introduction
    It offers a comprehensive structural analysis of the Generative AI domain within the context of fake news.
@@
    By clustering existing work, this survey identifies and categorizes the existing approaches.
@@
    It systematically enumerates the leading venues and research groups, providing insights into the epicenters of current research efforts.
@@
    This survey highlights unresolved issues and potential future research directions, thereby setting a roadmap for forthcoming investigations.
This survey is structured as follows: Section II examines existing surveys and underlining the need for this survey. Section III introduces the key concepts and aspects of the domain under consideration. Section IV details the methodological framework adopted for this study. Section V offers an in-depth review of recent research developments within the domain. Section VI delves into the evolution of the research topic, highlighting prominent venues. Section VII synthesizes the findings from the papers presented in this survey, distilling the essence of current research and its implications for the future.
@@ IV-D Selection Process
An initial search identified 95 articles. After removing duplicates, 86 articles were screened by title and abstract. This led to 31 articles being excluded for not meeting the inclusion criteria. A full-text review of the remaining articles resulted in 55 articles used for this survey.
@@ IV-F Source types
This survey categorizes sources into scientific articles, presentations, and books. Scientific articles, providing empirical evidence, form the empirical backbone of this analysis. Presentations, offering insights on current practices, enrich the discourse with real-world applications. Books, delivering comprehensive views, extend the understanding of AI’s societal impact.
@@ VI-A Evolution of the topic
Case study social media as most relevant distribution channel [30] [44], [34], [62], [70] [33], [61], [63] [60] [11]
Detection of Fake News [3] [31], [65], [67] [13], [18], [21], [5], [68], [32] [20], [66], [69], [12] [23] [22] [7]
Deepfakes as upcoming technology [25], [28], [26] [24], [27], [29], [14]
@@ VI-B Relevant venues
From the data collated, it is evident that the Association for Computational Linguistics and the AAAI Conference on Artificial Intelligence are prominent platforms for presenting seminal works, indicative of the strong reliance on advanced computational methods in this domain. The recurring presence of certain conferences, such as the International Conference on Pattern Recognition and the International Joint Conference on Artificial Intelligence, highlights the importance of AI and machine learning techniques in the development and assessment of fake news detection methodologies.
@@ VI-C Most active groups
The research landscape in generative AI, deepfakes, and fake news detection features contributions from a select number of academic groups and institutions globally. These entities have collectively propelled advancements in understanding and mitigating misinformation’s effects. Figure 3 illustrates the evolution and focus of research activities by leading groups from 2018 to 2023, underlining the growing efforts to combat disinformation and enhance generative AI technologies. Aside from this survey, only a few groups delve into generative AI within the context of fake news, indicating the field’s emerging nature. The involvement of multiple groups, rather than isolated researchers, adds significant weight to the research conducted in this domain.
Year2018201920202021202220232024NTU [71][19]LMU [72][19]Cam [73][56]AIT [42][47]PEC [74]
Figure 3: Timeline of research activities by leading groups from 2018 to 2023, illustrating the dynamic landscape of contributions in the fields of digital forensics, disinformation detection, and generative AI. Each colored box represents a significant milestone or publication by the respective research group.
@@ VI-C2 Other Research Entities
Focusing on cybersecurity and information manipulation, Pôle d’excellence cyber (PEC) provides insights into combating information manipulation through comprehensive expert reports. Their contributions, while not exclusively on generative AI, are vital for understanding cybersecurity aspects of fake news and deepfakes[74].
@@ VIII Conclusion
This survey has undertaken a systematic exploration of the interplay between Generative AI and fake news, filling a critical void in the existing corpus of literature. In Section II, this survey has substantiated our methodological approach, highlighting the essential nature of this exhaustive survey within the dynamic milieu of Generative AI and misinformation. By categorizing the existing body of work, this survey revealed diverse methodologies in the fight against fake news, facilitated by Generative AI. Section III established a foundational comprehension of the domain, shedding light on pivotal concepts and aspects crucial for understanding the broader ramifications of Generative AI in the realm of fake news. It presented a comprehensive structural analysis of the domain, identifying gaps that were previously unaddressed. Section IV delineated the methodological framework employed in this investigation, guaranteeing a methodical exploration.
@@
Section V delivered a comprehensive review of recent scholarly endeavors, accentuating the progress achieved in leveraging Generative AI for both the detection and creation of fake news. A systematic enumeration of leading venues and research groups showcased the focal points of current research, highlighting where significant advancements were being made. Section VI charted the developmental trajectory of this research area, spotlighting noteworthy venues and the contributions of eminent research collectives. The survey brought to light unresolved issues within the domain, suggesting avenues for future research and setting a clear direction for scholarly inquiry. It also spotlighted pressing research queries that propel further scholarly endeavors. Section VII amalgamated the insights derived from the survey, presenting a condensed perspective of the contemporary state of the art and delineating prospective avenues for domain advancement. By pinpointing the state of the art, it outlined the forefront of technological advancements in Generative AI, emphasizing its dual role in both generating and detecting fake news. These contributions collectively enhanced our understanding of the domain, providing a valuable resource for researchers, practitioners, and policymakers engaged in the ongoing battle against misinformation.
Section VI charted the trajectory of research on Generative AI and its interplay with fake news. Researchers initially focused on foundational Generative AI technologies. [57] and [17] provided the early groundwork. As Generative AI matured, the research pivoted towards its application in creating and detecting synthetic media. Significant contributions in this phase included [24] and [25]. Concurrently, the societal impact of fake news received attention, with studies like [30] and [33] examining its effects on public perception and social media.
@@
We hope that the identification of unresolved issues and future research directions will serve to mature and propel the field forward, emphasizing the critical need for innovative solutions in this problem that is highly relevant for our society. In that direction, our second focus on the state-of-the-art methodologies of Generative AI, for positioning it as a key tool in the ongoing fight against fake news.

### v2

@@ I Introduction
    Coming from a technological perspective, this work structures the Generative AI domain used and usable for Fake News.
@@
    From a research perspective, this work presents and clusters existing research streams and practical work and offers a categorization.
@@
    Also from a research perspective, this work structures the field by identifying the current publication hot spots (venues where papers were published).
@@
    In the same systematic, it identifies the research hotspots with the active groups.
  5. 5.
    Finally, this survey highlights open issues and extrapolates relevant future research directions, offering a research roadmap.
This survey is structured as follows: Section II provides the context of this research, showing that it is missing. Section III introduces the key concepts and aspects of Fake News and Generative AI. Section IV details the methodological applied for this work. Section V provides a comprehensive in-depth review of relevant recent research. Section VI dissects the evolution of the research topic and presents the current central publication venues and research groups. Section VII synthesizes the findings from the survey, distilling the essence of current research and its implications for the future.
@@ III The Domain
As Generative AI evolves, it simultaneously enhances the ability to create synthetic content and offers tools for its detection. This dual-use nature highlights the need for technologies that foster trust and authenticity. Sandner describes blockchain as the “digital notary of the 21st century—a guardian of truth in an age of uncertainty”[17]. Blockchain technology provides immutable proof of content authenticity and ensures transparency across decentralized networks, offering a promising solution to verify origins and detect tampering in digital content. This approach strengthens trust and accountability, helping counter misinformation and securing data integrity.
@@ IV-D Selection Process
An initial search identified 112 articles. After removing duplicates, 96 articles were screened by title and abstract. This led to 37 articles being excluded for not meeting the inclusion criteria. A full-text review of the remaining articles resulted in 59 articles used for this survey.
@@ IV-F Source types
This survey categorizes the sources into scientific articles, presentations, and books. Scientific articles, providing empirical evidence, form the empirical backbone of this analysis. Presentations, offering insights on current practices, enrich the discourse with real-world applications. Books, delivering comprehensive views, extend the understanding of AI’s societal impact.
@@ V-3 Case study social media as most relevant distribution channel
Public perceptions of deepfake technology can significantly influence its impact on society. For instance, [66] has illustrated that even a single deceptive video could instigate severe geopolitical tensions, highlighting the urgent need for robust verification measures.
Beyond content-focused strategies, examining social signals can further enhance the detection of misinformation. For example, [67] propose leveraging social noise and social entropy as measures of uncertainty, thus offering novel approaches to evaluate the credibility and trustworthiness of shared content.
@@ V-4 Detection of Fake News
While automated detection mechanisms are on the rise, human oversight remains crucial. As [73] demonstrate, computational approaches can sometimes outperform human detection rates, yet human involvement still plays a key role in the refinement and validation of detection tools.
@@ V-5 Deepfakes as upcoming technology
As shown by [74], older adults may face unique challenges in spotting deepfakes due to their differing identification strategies and experience with digital content. Moreover, demographic differences in detection abilities, such as those highlighted by [74], underscore the need for tailored training and resources for various user groups.
@@ VI-A Evolution of the topic
Case study social media as most relevant distribution channel [31] [45], [35], [63], [75] [34], [62], [64] [61] [11] [66], [67]
Detection of Fake News [3] [32], [68], [70] [13], [19], [22], [5], [71], [33] [21], [69], [72], [12] [24] [23] [7] [73]
Deepfakes as upcoming technology [26], [29], [27] [25], [28], [30], [14] [74]
@@ VI-B Relevant venues
From the data collated, it is evident that the Proceedings of the Association for Information Science and Technology, Association for Computational Linguistics and the AAAI Conference on Artificial Intelligence are prominent platforms for presenting seminal works, indicative of the strong reliance on advanced computational methods in this domain. The recurring presence of certain conferences, such as the International Conference on Pattern Recognition and the International Joint Conference on Artificial Intelligence, highlights the importance of AI and machine learning techniques in the development and assessment of Fake News detection methodologies.
@@
Journal Proceedings of the Association for Information Science and Technology [74], [66], [73], [67]
@@ VI-C Most active groups
The research landscape in generative AI, deepfakes, and Fake News detection features contributions from a select number of academic groups and institutions globally. These entities have collectively propelled advancements in understanding and mitigating misinformation’s effects. Figure 3 illustrates the evolution and focus of research activities by leading groups from 2018 to 2024, underlining the growing efforts to combat disinformation and enhance generative AI technologies. Aside from this survey, only a few groups delve into generative AI within the context of Fake News, indicating the field’s emerging nature. The involvement of multiple groups, rather than isolated researchers, adds significant weight to the research conducted in this domain.
Year2018201920202021202220232024NTU [76][20]LMU [77][20]Cam [78][57]AIT [43][48]PEC [79][80]
Figure 3: Timeline of research activities by leading groups from 2018 to 2024, illustrating the dynamic landscape of contributions in the fields of digital forensics, disinformation detection, and generative AI. Each colored box represents a significant milestone or publication by the respective research group.
@@ VI-C2 Other Research Entities
Focusing on cybersecurity and information manipulation, Pôle d’excellence cyber (PEC) provides insights into combating information manipulation through comprehensive expert reports. Their contributions, while not exclusively on generative AI, are vital for understanding cybersecurity aspects of Fake News and deepfakes[79, 80].
@@ VI-D1 Notable Gap in Literature: Generative AI and Fake News
VI-E The Next Frontiers in Fake News and Generative AI
Researchers initially focused on foundational Generative AI technologies. [58] and [18] provided the early groundwork. As Generative AI matured, the research became more and more focused on its application in creating and detecting synthetic media. Significant contributions in this phase included [25] and [26]. Concurrently, the societal impact of Fake News received attention, with studies like [31] and [34] examining its effects on public perception and social media.
@@ VIII Conclusion
This survey has undertaken a systematic exploration of the interplay between Generative AI and Fake News, filling a critical void in the existing corpus of literature. In Section II, this survey has substantiated our methodological approach, highlighting the essential nature of this exhaustive survey within the dynamic environment of Generative AI and misinformation. By categorizing the existing body of work, this survey revealed diverse methodologies in the fight against Fake News, facilitated by Generative AI. Section III established a foundational comprehension of the domain, shedding light on pivotal concepts and aspects crucial for understanding the broader ramifications of Generative AI in the realm of Fake News. It presented a comprehensive structural analysis of the domain, identifying gaps that were previously unaddressed. Section IV delineated the methodological framework employed in this investigation, guaranteeing a methodical exploration.
@@
Section V delivered a comprehensive review of recent scholarly endeavors, accentuating the progress achieved in leveraging Generative AI for both the detection and creation of Fake News. A systematic enumeration of leading venues and research groups showcased the focal points of current research, highlighting where significant advancements were being made. Section VI charted the developmental trajectory of this research area, spotlighting noteworthy venues and the contributions of eminent research collectives. The survey brought to light unresolved issues within the domain, suggesting avenues for future research and setting a clear direction for scholarly inquiry. It also spotlighted pressing research queries that propel further scholarly endeavors. Section VII amalgamated the insights derived from the survey, presenting a condensed perspective of the contemporary state of the art and delineating prospective avenues for domain advancement. By pinpointing the state-of-the-art, it outlined the forefront of technological advancements in Generative AI, emphasizing its dual role in both generating and detecting Fake News. These contributions collectively enhanced our understanding of the domain, providing a valuable resource for researchers, practitioners, and policy makers engaged in the ongoing battle against misinformation.
Section VI charted the trajectory of research on Generative AI and its interplay with Fake News. Over time, the focus of Generative AI research expanded from foundational model building to applying these models for both creation and detection of synthetic media. This gradual shift led to a broader awareness of the ethical, societal, and mitigation dimensions of Fake News, reflecting a transition from purely technical solutions to a more comprehensive perspective on hyper-realistic AI-generated content and its public impact. By early 2024, advanced models capable of producing highly realistic content across multiple formats highlighted the evolving sophistication of Generative AI systems. This progression underscored both the technological advancements achieved and the imperative for responsible governance to address associated challenges.
@@
The identification of unresolved issues and future research directions will serve to mature and propel the field forward. It emphasizes the critical need for innovative solutions for handling Fake News which is a key challenge for our society at this generation.